# Vision-Language Models for Edge Networks: A Comprehensive Survey

### v1

@@ III-B1 Data Selection and Pre-processing
Pre-processing data effectively is crucial for optimizing performance, particularly when dealing with heterogeneous data distributed across various edge devices. The process begins with Data Collection, where diverse data types such as images, text, and other relevant formats are gathered from multiple sources. The collected data must be segmented to ensure efficient processing across different environments, categorizing the data into Edge-Appropriate Data and Cloud-Appropriate Data. Edge-appropriate data generally consists of smaller, less complex datasets that can be processed in real-time on edge devices using lightweight models, while cloud-appropriate data involves more complex or sensitive information that necessitates extensive computational resources available in cloud environments [102, 103].
The next phase is Feature Extraction and Selection. During this step, relevant features are extracted from the raw data, enabling the child model on the edge device to process it efficiently. Feature selection determines which features should be processed locally and which should be sent to the cloud for further analysis, often using heuristics or lightweight models to assess data importance or complexity [104, 105].
To optimize further, Data Compression techniques minimize the bandwidth required for data transmission between edge and cloud. Standard methods include quantization, dimensionality reduction, and image compression. Local pre-processing on edge devices also helps reduce the volume of data transmitted, enhancing system efficiency [103, 104]. Advanced methodologies such as Asynchronous Aggregation and Cluster Pairing introduce an intermediate layer of edge servers between clients and the cloud, aggregating local models asynchronously to reduce communication overhead and speed up convergence. This method is effective in managing system heterogeneity [104]. Another approach is using Bioinspired Computing (BIC) algorithms, like Particle Swarm Optimization (PSO) and Genetic Algorithms, which address challenges in federated learning (FL), such as communication costs and system heterogeneity. These algorithms optimize resource allocation and data partitioning, ensuring relevant and manageable local processing [103]. Synchronous-Asynchronous Hybrid Update Strategy combines synchronous and asynchronous updates to mitigate staleness effects caused by Non-IID data, integrating local updates with global synchronization to enhance model accuracy and reduce idle times [104]. These pre-processing strategies are essential for enhancing the efficiency and performance of federated learning models in distributed, resource-constrained environments.
@@ III-B2 Model Choice on Edge and Cloud
Compression Techniques: Various model compression techniques enable deploying more complex models on resource-constrained edge devices by reducing model size and computational requirements without significantly compromising performance. Fig. 7 summarizes these techniques. These techniques include quantization, model compression, and knowledge distillation, among others. In Quantization, the precision of weights and activations is reduced from 32-bit floating-point to lower-bit representations (e.g., 16-bit or 8-bit integers), drastically reducing model size and computational overhead [110]. In Pruning, the process involves removing less significant neurons, weights, or layers to reduce size and complexity, achieving substantial reductions in model size and inference time [111]. on the other hand, Knowledge Distillation transfers knowledge from a larger, complex model (teacher) to a smaller model (student), enabling comparable performance with fewer parameters, useful for edge deployment [112].
On the other hand, other advanced techniques have emerged to further enhance model compression and efficiency. Neural Architecture Search (NAS) automates the design of efficient model architectures by searching a predefined space, optimizing for edge deployment [113]. Layer-wise Adaptive Rate Scaling (LARS) adjusts learning rates of different layers during training, combining with other compression methods to fine-tune performance [114]. Federated Dropout uses different subsets of a model’s parameters during training in a federated setting, reducing communication costs and yielding a smaller, more efficient model for edge deployment [115]. These techniques contribute to effective model deployment across edge and cloud environments, ensuring that models are well-suited to their respective operational constraints while maintaining high performance.
@@ III-B3 VLMs Implementation Across Distribution Locations
The setup of FL starts with the model partitioning, which determines the architecture of the parent model hosted on the cloud (more computation-heavy) and the child model hosted on the edge side (computationally lightweight). The Parent-Child Model Setup is important: The child model is local, light on resources, and only has to perform simple tasks; the Parent model is more complex and is responsible for pushing updates. A child model capable of processing independently for online inference with minimal on-device memory is achieved using the hierarchical architecture. On the other hand, the cloud model deals with model updates and computationally intensive tasks [116, 117].
Then, the next stage is delegating the jobs. As processes (real-time predictions, local data processing, etc.) are done in the edge model with low latency (or even with no latency), it stands to reason that the situation can be similar to Interaction with Cloud. On the other hand, the cloud model executes resource-intensive tasks, including model aggregation, complex analytics, and batch updates. This division of tasks permits leveraging the available edge and cloud resources efficiently [118, 119].
@@
Once models are separated, a data distribution policy is implemented. In contrast, local data handling is seen as relevant to processes on the edge, which generate updates or predictions locally and communicate less frequently with the cloud.
@@
    Selective Uploading: These methods determine which part of the processed data are essential to upload to the cloud, with only significant information related to the update will be sent or data needed deeper analysis [120, 121]. Keeping only the most relevant data to upload reduces the amount of data transferred over the network and increases the system’s efficiency.
@@
    Model Update Mechanism: It maintains the synchronization of the edge model with the compatibility and component updates of the cloud model. Local training: The child model can be trained using edge devices where the input data is collected locally, leaving the limited computational capacity behind.
@@
    Federated Averaging Algorithm: The updates from the different edge devices are averaged on the cloud to update the global model to achieve better hypotheses. Federated Averaging enables the international model to gain insights from non-independent data sources without revealing them to others, which minimizes the probability of data leakages [118, 115].
@@
    Model Synchronization: This ensures that the child model stays frozen with global model updates. Periodic sync lets the cloud return model parameters to the edge device so the child model can benefit from collective learning happening on devices. In this way, the model remains accurate and efficient across the FL system while simultaneously synchronizing and performing local training [119, 122].
@@ III-B4 Post Processing and Evaluation
Post-processing and evaluation are essential stages in the FL system life-cycle, ensuring that both local and global models achieve optimal performance across heterogeneous environments. The first step in post-processing is Local Evaluation, which continues to measure the child model performance on the edge device using local data. It is important to notice performance degradation, which signals a model to be updated or updated from the cloud. New approaches utilize continual learning and on-the-fly performance tracking to adjust the model dynamically to the fluctuating edge contexts [123]. As indicated by [124], lightweight performance estimation and anomaly detection algorithms are deployed locally to monitor changes in the model behavior, triggering invocation of deployment requests to the cloud.
The Feedback Loop is an integral part of local evaluation as it helps understand how the edge model performs and what kind of data or scenarios the edge model finds difficult. This feedback allows the global model to be refined by showing where more training or tweaks are needed. Many researchers use reinforcement learning and meta-learning to let the edge model self-improve in the long run. At the same time, this data also enhances the global model [125].
Global Model Evaluation: It merges data and updates from multiple edge devices in the cloud to comprehensively evaluate the global model. This aggregation enables the cloud to evaluate global model performance in heterogeneous environments and user space. Federated evaluation frameworks have been developed using privacy-preserving methods to aggregate performance metrics without revealing individual user data [126]. These frameworks leverage secure aggregation and differential privacy to ensure correct and secure global model evaluation.
Model Tuning on the Aggregation: A global model is fine-tuned on the server side based on the aggregated updates from edge devices. This tuning process captures edge-level nuances in data dynamics and contributes to increasing the generalizability of the global model. Sophisticated approaches, such as federated hyperparameter tuning and federated Bayesian optimization, are employed to ensure the global model remains adapted to different conditions and devices [127].
@@ III-B5 Deployment and Continuous Learning
Deployment and continuous learning are essential components of FL systems, ensuring that models are effectively updated, scaled, and adapted to changing environments. However, recent progress has led to various state-of-the-art (SOTA) methods that improve these processes to enable more robust and scalable FL.
Model Update Deployment — Frequent model parameters are released from the cloud to edge devices as model updates to ensure Ongoing Learning of the Edge devices. This integration ensures that edge devices are always using the latest model updates. Emerging approaches focus on transferring lightweight model updates instead of complete model files and applying differential synchronization mechanisms to limit the pass-through data between cloud and edge, mitigating latency and bandwidth consumption [128]. Models with sparse updates, such as federated dropout, only require the update of critical model parameters, leading to higher efficiency in using limited network resources, as proposed by the work in [129].
Edge Device Management ensures efficient deployment of updates across multiple edge devices. Modern approaches leverage orchestration frameworks that automate deployment, ensuring timely updates for all devices. These frameworks often employ decentralized deployment strategies, where updates are propagated in a peer-to-peer fashion, reducing load on central servers and improving scalability [130]. Adaptive deployment techniques manage heterogeneous edge environments, tailoring updates to each device’s capabilities [131].
A Scalable Architecture is essential for handling an increasing number of edge devices. SOTA methods focus on creating architectures that can scale horizontally by adding more edge nodes and vertically by enhancing cloud resources. Cloud-native technologies like Kubernetes and containerization have been widely adopted to manage the deployment of FL models across large clusters of edge devices [132]. These technologies allow for dynamic scaling, ensuring that the system can handle the aggregation of updates from a growing number of devices without bottlenecks.
Adaptation to New Data is another critical aspect of continuous learning. As data environments evolve, FL systems must adapt to new data types and patterns to maintain model accuracy. Techniques such as continual learning and federated meta-learning enable models to adapt without forgetting previously learned information [133]. Moreover, integrating reinforcement learning into FL systems allows models to adjust learning strategies dynamically based on environmental changes, ensuring sustained performance even in non-stationary environments [134].
Existing Vision-Language Models face substantial limitations when deployed on edge devices due to high computational and memory requirements for complex visual and language processing. These models are often designed for cloud environments with abundant resources, making them challenging to run efficiently on resource-constrained devices such as IoT gateways and mobile robots. Edge devices typically lack the hardware for large-scale data processing, leading to issues with latency, limited real-time processing capabilities, and restricted power efficiency. Additionally, current models struggle to adapt across different visual data modalities (e.g., RGB and depth images) and frequently rely on centralized cloud-based data aggregation, raising concerns about data privacy and network dependency.
@@
To overcome these limitations, it is essential to develop Vision-Language Models optimized for edge deployment. These models should incorporate lightweight architectures that balance computational efficiency with high performance. Techniques such as quantization, pruning, and knowledge distillation effectively reduce model size and computational demand without compromising accuracy. Advanced methodologies like Neural Architecture Search and Federated Learning enhance adaptability in distributed, resource-limited environments while maintaining data privacy [135]. Future models must also support flexible, real-time processing across various visual modalities and incorporate continual learning mechanisms to adapt dynamically to new data patterns and evolving tasks in diverse edge applications.
@@ IV Recent Advances
Applications of lightweight VLMs are expanding rapidly across various industries, driven by the need for efficient, on-device multi-modal processing. There are many tasks where VLMs can help, including Text-Image Retrieval, image captioning, Question and answer Classification, object detection, and segmentation, as shown in Figure 8. These models are being deployed in fields such as autonomous systems, healthcare, surveillance, environmental monitoring, and many other applications. Table IV summarizes the applications we covered in this survey.
@@
IV-A VLM in healthcare
IV-B VLMs in Environmental Monitoring and Aerial Imaging
@@ V Open Challenges
This section presents novel challenges in advancing VLMs for edge applications. Recent surveys highlighted several critical issues. Table IV shows the challenges we introduce and the difference between these and previous discussions. The alignment gap between visual and textual modalities reduces effectiveness in tasks like image captioning and question answering [34, 35]. VLMs depend on large-scale datasets, which are costly to develop and limit accessibility for smaller institutions [37]. High computational demands hinder deployment on resource-constrained devices [41, 36], and VLMs often fail to generalize across domains, underperforming on novel tasks [38]. Additionally, parameter inefficiencies in dual-stream architectures increase memory usage, making them less suitable for real-time or edge-based applications [34]. These challenges underscore the need for innovative approaches distinct from current solutions in the literature.
The surveyed papers address various challenges associated with deploying LLMs on edge devices. Cai et al. [158] discuss high computational demands and reliance on large-scale datasets, limiting VLM deployment on resource-constrained devices, along with an alignment gap between visual and textual modalities. Bhardwaj et al. [159] emphasize optimization to balance computational demands, energy efficiency, and model scalability for edge devices. Lu et al. [160] address generalizing VLMs for edge AI, including AI integration into wearables, hardware miniaturization, and user-friendly interfaces. Qu et al. [161] propose mobile edge intelligence to bridge cloud and on-device AI, highlighting privacy concerns, latency issues, and resource limitations of edge devices. Lin et al. [162] focus on deploying VLMs in 6G edge networks, identifying challenges like response times, bandwidth costs, and privacy risks. Yuan et al. [163] discuss energy efficiency and computational constraints for mobile devices running VLM inference tasks, proposing approaches to enhance energy efficiency. Qu et al. [164] examine limitations of on-device LLMs due to edge devices’ constrained capacity, advocating mobile edge intelligence to reduce latency and privacy issues. Chen et al. [165] explore LLM integration into edge intelligence, focusing on adaptive applications and throughput challenges for small models on edge devices. Lee et al. [166] discuss adapting Vision Transformers for mobile and edge devices, emphasizing computational efficiency and implementation challenges on resource-limited devices.
@@ V-A Compressed Light Weight VLMs for Edge Networks
How does one propose novel compressed light-weight VLMs for edge networks with a reasonable performance? Generally, it is challenging to extend VLMs to the network edge due to high computing resource requirements for training complex VLM models with large number of parameters. To do resolve this challenge, one way is to use compression schemes. Several model compression techniques are being explored to run large VLMs on edge devices. Knowledge distillation and quantization are two prominent approaches to reduce model size and minimize computation time without sacrificing too much performance. Knowledge distillation transfers knowledge from a larger model to a smaller one, allowing the lightweight version to retain the essential functionalities of the original. For instance, the EdgeVL framework utilizes dual-modality knowledge distillation to support both RGB and non-RGB images, reducing the model size by up to 93 times and improving accuracy by 15% on edge devices [156, 157]. Similarly, quantization-aware training helps adapt models for lower precision, conserving memory and power while maintaining accuracy. Although these existing schemes performs better, we might need novel compression schemes for specific applications (e.g., VLMs for medical imaging at the network edge and remote sensing assisted by drones) on VLMs at the network edge. Therefore, there is a need for more novel schemes for various applications. For instance, MiniVLM utilizes knowledge distillation to significantly reduce the model size for edge deployment while maintaining over 90% of the original performance, making it suitable for cross-modal retrieval tasks [82]. Another approach, DIME-FM, distills large VLMs like CLIP into smaller, more efficient models using unpaired image and text data. This method maintains transferability and robustness, making it suitable for resource-constrained edge applications such as real-time image and text matching tasks [167].

### v2

@@ III-B1 Data Selection and Pre-processing
Effective pre-processing is crucial for optimizing performance, especially with heterogeneous data across edge devices. It begins with Data Collection, where various data types (images, text, etc.) are gathered from multiple sources. Data is then segmented into Edge-Appropriate and Cloud-Appropriate categories. Edge-appropriate data is simpler and can be processed in real-time on devices using lightweight models, while cloud-appropriate data requires more complex processing and resources [94, 95].
In Feature Extraction and Selection, relevant features are derived from raw data to support efficient processing by edge models. Feature selection helps determine which features are processed locally and which are sent to the cloud, often using heuristics or lightweight models [96, 97].
Data Compression further reduces bandwidth for edge-cloud transmission using methods like quantization, dimensionality reduction, and image compression. Local pre-processing helps minimize transmitted data volume [95, 96]. Advanced methods like Asynchronous Aggregation and Cluster Pairing introduce intermediate edge servers to aggregate local models asynchronously, reducing communication overhead and accelerating convergence [96]. Bioinspired Computing (BIC) algorithms, such as PSO and Genetic Algorithms, help tackle communication costs and system heterogeneity by optimizing resource allocation and data partitioning [95]. The Synchronous-Asynchronous Hybrid Update Strategy mitigates staleness from Non-IID data by combining local updates with global synchronization, improving accuracy and reducing idle time [96]. These pre-processing strategies are vital for efficient federated learning in distributed, resource-constrained settings.
@@ III-B2 Model Choice on Edge and Cloud
Compression Techniques: Model compression enables deploying complex models on resource-constrained edge devices by reducing size and computational demands without major performance loss. Fig. 7 summarizes key techniques, including quantization, pruning, and knowledge distillation. In Quantization, weights and activations are reduced from 32-bit floating-point to lower-bit formats (e.g., 16-bit or 8-bit integers), significantly reducing size and computation [102]. Pruning removes less significant neurons, weights, or layers, cutting model size and inference time [103]. Knowledge Distillation transfers knowledge from a large teacher model to a smaller student model, achieving comparable performance with fewer parameters, making it ideal for edge deployment [104].
Other advanced methods further improve compression and efficiency. Neural Architecture Search (NAS) automates the design of efficient architectures by exploring a search space optimized for edge use [105]. Layer-wise Adaptive Rate Scaling (LARS) adjusts learning rates across layers during training, and can be combined with other methods to fine-tune performance [106]. Federated Dropout uses different subsets of a model’s parameters during training in a federated setup, reducing communication costs and yielding compact, efficient models for edge deployment [107]. These methods support effective deployment across edge and cloud environments, balancing performance with operational constraints.
@@
Balancing Model Compression and Accuracy
Implementing model compression techniques such as pruning, quantization, and knowledge distillation involves a trade-off between computational efficiency and accuracy. Aggressive compression reduces model size and inference time but can degrade performance. To balance this, the following strategies are commonly employed:
  • Incremental Compression: Gradually applying compression enables accuracy monitoring at each step, helping find an optimal trade-off. [108] showed that structured pruning applied incrementally achieved substantial size reduction with minimal accuracy loss.
  • Layer-wise Sensitivity Analysis: Assessing layer sensitivity guides which layers can handle more aggressive compression. Less sensitive layers can be pruned or quantized more without major performance loss. [109] demonstrated that this approach preserves performance while lowering complexity.
  • Hybrid Compression Techniques: Combining methods like pruning, quantization, and distillation leverages their strengths. Ref.  [110] showed that such combinations reduce size effectively without compromising accuracy.
  • Adaptive and Dynamic Compression: Adaptive methods like dynamic quantization adjust precision based on data complexity. Ref. [111] proposed a joint distillation and quantization method that dynamically adjusts parameters for efficient compression.
  • Domain-specific Performance Metrics: Defining accuracy thresholds based on application needs ensures compression does not exceed acceptable degradation. Ref. [108] highlighted aligning strategies with domain-specific metrics for practical effectiveness.
Implementing these strategies requires careful consideration of the specific application and deployment environment to maintain a balance between efficiency and accuracy.
@@ III-B3 VLMs Implementation Across Distribution Locations
The setup of FL begins with model partitioning, which defines the architecture of the parent model on the cloud (computation-heavy) and the child model on the edge (lightweight). The Parent-Child Model Setup is key: the child model runs locally with minimal resources, handling simple tasks; the parent model is more complex and handles updates. A hierarchical architecture enables the child model to perform online inference independently with minimal on-device memory, while the cloud model manages updates and intensive computation [112, 113].
The next stage involves delegating tasks. Edge models handle real-time predictions and local processing with low or no latency, while the cloud model manages resource-intensive tasks such as model aggregation, complex analytics, and batch updates. This division allows efficient use of both edge and cloud resources [114, 115].
@@
Once models are separated, a data distribution policy is applied. Local data handling on the edge supports generating updates or predictions with minimal cloud interaction.
@@
    Selective Uploading: Determines which processed data are essential for cloud upload, sending only significant updates or data requiring deeper analysis [116, 117]. This reduces network usage and enhances system efficiency.
@@
    Model Update Mechanism: Maintains synchronization between the edge model and cloud model updates. Local training allows the child model to learn from on-device data without exceeding limited computation capacity.
@@
    Federated Averaging Algorithm: Updates from edge devices are averaged on the cloud to refine the global model, enabling learning from non-independent data sources while minimizing data leakage risk [114, 107].
@@
    Model Synchronization: Ensures the child model incorporates global updates via periodic syncing. The cloud returns updated parameters to edge devices, enabling local models to benefit from collective learning and maintain accuracy [115, 118].
Privacy and Security Considerations
Deploying VLMs on edge devices necessitates robust privacy and security measures to protect sensitive data. Several advanced techniques have been developed to address these concerns:
Differential Privacy (DP): DP introduces controlled noise to the data or model updates, ensuring that individual data points cannot be distinguished, thereby preserving privacy. In federated learning scenarios, DP has been effectively applied to protect client data during model training [119].
Secure Aggregation: This technique ensures that the server can aggregate model updates from clients without accessing individual contributions. Methods utilizing multiparty homomorphic encryption have been proposed to achieve efficient and secure aggregation in federated learning [120].
Homomorphic Encryption (HE): HE allows computations to be performed directly on encrypted data without decryption, maintaining data confidentiality throughout the process. Implementations like OpenFHE have been developed to support such privacy-preserving computations [121].
Integrating these techniques into the deployment of VLMs on edge devices enhances data privacy and security, addressing potential vulnerabilities in decentralized networks.
@@ III-B4 Post Processing and Evaluation
Post-processing and evaluation are critical in the FL system life-cycle to ensure both local and global models perform well across heterogeneous environments. The first step is Local Evaluation, which assesses child model performance on edge devices using local data. Detecting performance degradation indicates when updates from the cloud are needed. Recent approaches use continual learning and on-the-fly performance tracking to adapt models dynamically to changing edge contexts [122]. As noted by [123], lightweight performance estimation and anomaly detection run locally to monitor behavior and trigger cloud deployment requests.
The feedback loop is integral to local evaluation, helping identify difficult scenarios or data for the edge model. This feedback improves the global model by highlighting areas needing more training. Reinforcement learning and meta-learning techniques are often used to help edge models self-improve over time, also benefiting the global model [124].
Global Model Evaluation aggregates updates and data from multiple edge devices in the cloud, allowing a comprehensive assessment across varied environments and user contexts. Federated evaluation frameworks use privacy-preserving techniques to gather performance metrics without exposing user data [125]. Secure aggregation and differential privacy are key to enabling accurate and private evaluation.
Model Tuning on the Aggregation: The global model is fine-tuned on the server using updates from edge devices, capturing local data nuances and improving generalizability. Advanced strategies like federated hyperparameter tuning and federated Bayesian optimization ensure the global model remains effective across devices and conditions [126].
@@ III-B5 Deployment and Continuous Learning
Deployment and continuous learning ensure FL models are updated, scaled, and adapted to evolving environments. Recent advances have introduced SOTA methods to enhance robustness and scalability.
Model Update Deployment — Model parameters are frequently pushed from the cloud to edge devices to support ongoing learning. New approaches focus on lightweight updates and differential synchronization to reduce data exchange, minimizing latency and bandwidth use [127]. Sparse update techniques like federated dropout send only essential parameters, improving efficiency [128].
Edge Device Management automates update deployment across devices using orchestration frameworks. Decentralized strategies (e.g., peer-to-peer) reduce server load and improve scalability [129]. Adaptive techniques tailor updates to each device’s capabilities [130].
A Scalable Architecture is key for growing edge networks. SOTA solutions scale horizontally by adding edge nodes and vertically via cloud resource expansion. Cloud-native tools like Kubernetes and containerization enable dynamic scaling across large device clusters [131].
Adaptation to New Data is vital for maintaining accuracy as data evolves. Continual learning and federated meta-learning help models adapt without forgetting past knowledge [132]. Reinforcement learning allows dynamic strategy adjustments in non-stationary environments [133].
Current Vision-Language Models are difficult to deploy on edge devices due to high resource requirements. Designed for cloud use, they face challenges like latency, poor real-time performance, and power inefficiency. They also struggle with diverse visual modalities and rely on centralized data aggregation, raising privacy and dependency concerns.
@@
To address this, VLMs for edge should use lightweight architectures that balance performance and efficiency. Techniques like quantization, pruning, and knowledge distillation reduce size and computation without losing accuracy. Neural Architecture Search and Federated Learning improve adaptability and privacy in distributed settings [134]. Future models must support real-time, multimodal processing and continual learning for dynamic edge environments.
@@ IV Recent Advances
Applications of lightweight VLMs are expanding rapidly across various industries, driven by the need for efficient, on-device multi-modal processing. There are many tasks where VLMs can help, including Text-Image Retrieval, image captioning, Question and answer Classification, object detection, and segmentation, as shown in Figure 8. These models are being deployed in autonomous systems, healthcare, surveillance, environmental monitoring, and many other applications. Table IV-B summarizes the applications we covered in this survey.
IV-A VLM in healthcare
IV-B VLMs in Environmental Monitoring and Aerial Imaging
@@ V Open Challenges
This section presents novel challenges in advancing VLMs for edge applications. Recent surveys have identified several critical issues. Table IV highlights the challenges we introduce and how they differ from prior discussions. The alignment gap between visual and textual modalities limits performance in tasks like image captioning and question answering [34, 35]. VLMs also rely on large-scale datasets, which are expensive to develop and restrict access for smaller institutions [37]. High computational demands make deployment difficult on resource-constrained devices [41, 36], and VLMs often struggle to generalize across domains, underperforming on novel tasks [38]. Parameter inefficiencies in dual-stream architectures further increase memory usage, making them less suitable for real-time or edge-based applications [34]. These issues call for innovative solutions.
Surveyed papers explore several challenges in deploying LLMs on edge devices. Cai et al. [87] highlight computational demands and dataset reliance, along with the visual-textual alignment gap. Bhardwaj et al. [157] focus on optimizing computational load, energy efficiency, and model scalability. Lu et al. [158] discuss generalizing VLMs for edge AI, covering wearables, hardware miniaturization, and usability. Qu et al. [159] propose mobile edge intelligence to bridge cloud and device-side AI, addressing privacy, latency, and resource constraints. Lin et al. [160] examine VLM deployment in 6G edge networks, citing response time, bandwidth, and privacy challenges. Yuan et al. [161] emphasize energy-efficient inference on mobile devices. Qu et al. [162] critique on-device LLMs due to limited capacity, advocating edge intelligence to reduce latency and enhance privacy. Chen et al. [163] explore LLMs in edge intelligence, focusing on adaptive use and throughput challenges for small models. Lee et al. [164] address adapting Vision Transformers for edge use, emphasizing efficiency and deployment challenges on constrained devices.
@@ V-A Compressed Light Weight VLMs for Edge Networks
How does one propose novel compressed light-weight VLMs for edge networks with a reasonable performance? Deploying VLMs at the edge remains challenging due to the high computational demands of training large models with many parameters. A common solution is model compression. Techniques such as knowledge distillation and quantization are widely used to reduce model size and computation without significant performance loss. Knowledge distillation transfers knowledge from a large model to a smaller one, allowing it to retain essential functionalities. For example, the EdgeVL framework employs dual-modality distillation to support RGB and non-RGB images, reducing model size by up to 93× and improving accuracy by 15% on edge devices [155, 156]. Quantization-aware training further adapts models to lower-precision formats, saving memory and power while preserving accuracy. Although existing methods perform well, many edge scenarios—such as medical imaging or drone-based remote sensing—require specialized compression schemes. MiniVLM, for instance, applies knowledge distillation to achieve substantial model reduction while maintaining over 90% of the original performance, making it suitable for cross-modal retrieval tasks [61]. Another method, DIME-FM, distills large models like CLIP using unpaired image-text data to produce compact, robust models suited for tasks like real-time image-text matching in resource-constrained environments [165]. These examples highlight the ongoing need for tailored lightweight VLMs optimized for diverse edge applications.